A quantitative data set has mean 23 and standard deviation 1. At least what percentage of the observations lie between 21 and 257 A. 68% B. 95% С. 25 00% OD, 75.00%
Answer #1

The correct answer is 75.00%

So correct option is D

Because by Chebyshev's inequality, we get the value of k such that

k = rac{x -mu }{sigma }

plug x = 25,  mu = 23' and  sigma = 1; in the above equation.

25 - 23

Then we get at least probability as ( 1-(1/k^2) ) = 1 - (1/2^2) = 1 = (1/4) = 3/4 = 0.75

Multiply it by 100 , so we get the answer in percentage form.

That is 0.75*100 = 75.00%